OpenStudy (anonymous):

2. 3x+2y=7 -3+11=y a.(6,-3 b.(6,-7) c(-4,9/12) d.(5,-4)

OpenStudy (anonymous):

i think the answer is B

OpenStudy (a1234):

Do you have something missing in the second equation?

OpenStudy (anonymous):

yea its -3x im sorry

OpenStudy (ankit042):

K so add the two eauations to proceed you will form a equation in Y

OpenStudy (anonymous):

b is not the answer @Lovecraze1601

OpenStudy (a1234):

3x+2y=7 -3x+11=y I'll use substitution: 3x + 2(-3x + 11) = 7 3x - 6x + 22 = 7 -3x = -15 x = 5 3(5) + 2y = 7 15 + 2y = 7 2y = -8 y = -4 (5, -4)

OpenStudy (anonymous):

so thats how u do it

OpenStudy (a1234):

Yup :).

OpenStudy (anonymous):

can u help we with 1 more

OpenStudy (a1234):

Sure.

OpenStudy (anonymous):

using substitution 2x-y=-7 4x-y=-4 a.(-1.5,4) b.(1.5,10) c.(4,-1.5) d.(-1.5,-2)

OpenStudy (a1234):

Solve for y from any of the two equations: 2x - y = -7 -y = -2x - 7 4x-y=-4 4x - 2x - 7 = -4 2x = 3 x = 1.5 4(1.5) - y = -4 6 - y = -4 -y = -10 or y = 10 (1.5, 10)

OpenStudy (anonymous):

thank you sooooooooooooo much

OpenStudy (a1234):

You're welcome!

OpenStudy (anonymous):